### v1

# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

Abstract
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. Achieving such reconstruction with limited training data requires generalization beyond the training set, a task known as zero-shot prediction. Despite its importance, we still lack theoretical guidelines for achieving efficient and accurate reconstruction. In this paper, we provide a theoretical analysis of two widely used models for translating brain activity to latent image features. We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize how each model behaves across data scales. We first show that the naive linear regression model, which uses a shared set of input variables for all outputs, suffers from “output dimension collapse” at small data scales, restricting generalization beyond the training data. We then mathematically characterize the prediction error of the sparse linear regression model by deriving formulas linking prediction error with data scale and other problem parameters. Leveraging the sparsity of the brain-to-feature mapping, this approach enables accurate zero-shot prediction even at small data scales without trapping in output dimension collapse. Our results provide a theoretical guideline for achieving zero-shot reconstruction and highlight the benefits of variable selection in brain-to-image reconstruction.
keywords:
Brain decoding , Brain-to-image reconstruction , Zero-shot prediction , Sparse linear regression
[…]
1 Introduction
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. To uncover neural representations and move toward practical applications in medicine and industry, it will be necessary to reconstruct a broad range of subjective visual experiences. Nevertheless, a fundamental constraint governs this task: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training. To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, namely zero-shot prediction (Shirakawa et al., 2025). Despite its importance, we still lack theoretical guidelines for achieving accurate reconstruction under these constraints.
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). A translator module converts brain activity into latent features that represent the image content, and a generator module transforms these features into reconstructed images. This paper focuses on the translator module, trained on a limited pairs of brain activity and latent features. We quantify this limitation by the data scale, the ratio of the number of training pairs to the dimensionality of the latent features. To realize zero-shot reconstruction, the translator must be able to predict valid latent features for unseen stimuli at small data scales.
Figure 1: Translator–Generator pipeline. The translator converts brain activity into latent features, and the generator transforms latent features into reconstructed images. A naive multivariate linear regression model and a sparse multivariate linear regression model are commonly used as translators. Whereas a naive model uses a single shared set of input variables to predict all output dimensions, a sparse model allows each output dimension to select its own subset of input variables.
Two linear regression models are widely used as translators: naive multivariate linear regression and sparse multivariate linear regression. A naive multivariate linear regression model, exemplified by ridge regression, uses a single shared set of input variables to predict all output dimensions and has been adopted in several high-profile studies (Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This strategy is straightforward and low-cost, yet Shirakawa et al. (2025) reports that such models fail to generalize beyond the training domain. They attribute this failure to “output dimension collapse,” in which predictions are confined to the span of the training outputs under dataset bias. In contrast, a sparse multivariate linear regression model allows each output dimension to select its own subset of input variables. Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
In this paper, we provide a theoretical analysis of translator prediction under small data scales. First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales cause output dimension collapse in naive multivariate linear regression. Using the mathematically derived best prediction under this collapse, we evaluate its impact on reconstruction pipelines that employ naive translators. Second, we examine how sparsity can enable accurate prediction at small data scales by characterizing the prediction error of naive and sparse translators. Our contributions are as follows:
  • 1.
    We show that reconstruction pipelines using naive linear regression suffer from output dimension collapse at small data scales.
  • 2.
    We mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters.
  • 3.
    We demonstrate that sparse linear regression can escape output dimension collapse and achieve accurate prediction at small data scales by leveraging sparse brain-to-feature structure.
2 Preliminaries
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). In this framework, a module called the translator converts brain activity into latent features, and a module called the generator transforms these features into reconstructed images. The translator is typically a linear regression model trained on paired brain activity and latent feature data. The generator is typically a pre-trained or pre-defined model that generates images from the predicted features. Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features (e.g., CNN, CLIP).
In this study, we focus on the translator module, which maps brain activity $\bm{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ to latent features $\bm{y}\in\mathbb{R}^{d_{\mathrm{out}}}$. The training set consists of $n$ paired samples of brain activity and latent features, denoted by $\mathcal{D}=\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The latent features $\bm{y}$ are usually high-dimensional representations because they represent a broad spectrum of visual stimuli. Meanwhile, the number of available brain activity samples $n$ is limited; $n$ is much smaller than $d_{\mathrm{out}}$. We therefore define the data scale as the ratio ${n}/{d_{\mathrm{out}}}$, which is typically much smaller than 1 in reconstruction studies. To achieve zero-shot reconstruction, the translator must predict valid latent features for unseen stimuli, even under small data scales.
A common approach to implementing a translator is naive multivariate linear regression, which utilizes a single, shared set of input variables to predict all output dimensions. Representative methods include ridge regression, kernel linear regression, and partial least squares (PLS). These methods differ in how they regularize or project the data, but still rely on shared inputs for all output dimensions.
In contrast, sparse multivariate linear regression allows each output value to select its subset of input variables. According to the classic taxonomy (Guyon and Elisseeff, 2003), selection strategies fall into three groups: wrapper methods (Kohavi and John, 1997), which iteratively search feature subsets by maximizing predictive performance; embedded methods (e.g., Lasso (Tibshirani, 1996), Automatic Relevance Determination (ARD) (MacKay, 1992)), which embed sparsity directly in the training objective; and filter methods (e.g., CFS (Hall, 2000), SIS (Fan and Lv, 2008) ), which rank inputs independently of the learner using simple scores (e.g., correlation, mutual information). In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and in practice, the runtime cost scales as wrapper $\gg$ embedded $\gg$ filter.
Previous theoretical work demonstrates that sparsity is particularly beneficial when the sample size is smaller than the dimensionality of the data. For a truly $s$-sparse input-output mapping, information-theoretic bounds show that an ideal estimator can attain near-oracle prediction error once $n\gtrsim s\log(d_{\mathrm{in}}/s)$ (Wainwright, 2009a). The Lasso is the most thoroughly analyzed among practical sparse methods, and theory confirms it reaches essentially the same rate and achieves accurate prediction with $n=\mathcal{O}(s\log(d_{\mathrm{in}}))$ (Donoho, 2006; Wainwright, 2009b). The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ is exponentially large in $n$ (Fan and Lv, 2008).
3 Output dimension collapse by small data scale
Although both sparse and naive (non-sparse) models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This section focuses on the naive translator and shows a critical limitation at small data scales.
In brain-to-image reconstruction, the translator must map brain activity $\bm{x}$ onto a high-dimensional image feature space $\bm{y}$ while relying on a limited training sample. If we use a naive multivariate linear regression under such scenarios, a phenomenon “output dimension collapse” (ODC) can occur, whereby the predictions become restricted to the subspace determined by the training outputs (Shirakawa et al., 2025). This issue is especially critical for zero-shot prediction when the training subspace is low-dimensional, because the model inevitably compresses unfamiliar features into that subspace.
To understand why this collapse occurs, let us examine the predicted features of ridge regression. First, we learn a weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ by minimizing the sum of squared errors plus a regularization term:
$\hat{W}$ $=\underset{W}{\operatorname{argmin}}\sum_{i=1}^{n}\lVert\bm{y}^{(i)}-W^{\top}\bm{x}^{(i)}\rVert^{2}+\lambda\lVert W\rVert^{2}_{F}$ (1)
$=\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}{X_{\mathrm{tr}}}^{\top}{Y_{\mathrm{tr}}}.$ (2)
Here, $X_{\mathrm{tr}}=[\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(n)}]^{\top}\in\mathbb{R}^{n\times d_{\mathrm{in}}}$ and $Y_{\mathrm{tr}}=[\bm{y}^{(1)},\bm{y}^{(2)},\ldots,\bm{y}^{(n)}]^{\top}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ are matrices which arrange the brain activity data and latent features of the training data row-wise, respectively. For a test input brain activity $\bm{x}_{\mathrm{te}}$, the prediction is
$\hat{\bm{y}}_{\mathrm{te}}$ $={\hat{W}}^{\top}\bm{x}_{\mathrm{te}}$ (3)
$={Y_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}$ (4)
$=Y_{\mathrm{tr}}^{\top}\bm{m}=\sum_{i=1}^{n}m_{i}\bm{y}^{(i)}$ (5)
where $\bm{m}={X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}\in\mathbb{R}^{n}$, and $m_{i}\in\mathbb{R}$ is the $i$-th element of $\bm{m}$. Clearly, the prediction $\hat{\bm{y}}_{\mathrm{te}}$ lies in the subspace spanned by the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. In other words, no matter what the test input is, the predicted output is constrained to that subspace (Figure 2A). Importantly, this property is not the familiar one that confines predictions to the span of the training inputs $\left\{\bm{x}^{(i)}\right\}_{i=1}^{n}$, but instead that it confines predictions to the span of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. This subspace is determined solely by the training outputs and does not depend on the input brain data; given the same training outputs, the same property holds regardless of measurement modality (e.g., fMRI, MEG), subject, or decoding task (e.g., perception, imagery, or illusion). Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see A.1 for details).
Figure 2: Output dimension collapse. (A) The predictions become restricted to a low-dimensional subspace determined by the training outputs. (B) The best prediction within the training feature subspace provides the lower bound for prediction error in naive multivariate linear regression.
A key condition that leads to output dimension collapse is when the training outputs lie in a low-dimensional subspace. Shirakawa et al. (2025) showed that such a collapse occurs when the training data lack sufficient diversity and the latent features exhibit a cluster structure. In such scenarios, the predicted outputs are pulled toward the clusters in the training data, making generalization to unseen stimuli difficult.
In contrast, we focus here on a different but equally critical factor: small data scales. Even in the absence of explicit cluster structure, having $n<d_{\mathrm{out}}$ inherently forces all predictions into an at-most $n$-dimensional subspace. As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe output dimension collapse. Consequently, regardless of whether the training outputs form visible clusters, small data scales alone can lead to the same fundamental difficulty in zero-shot prediction.
From another perspective, output dimension collapse due to small data scale is an inevitable consequence of reducing the number of free parameters in the model. The model must learn a weight matrix with many parameters to predict high-dimensional latent features, but the available training samples are insufficient to identify all of them. Therefore, the model must reduce the effective number of free parameters by imposing additional constraints. Naive linear regression accomplishes this by confining all predictions to the linear span of the training outputs. This implicit constraint is precisely what leads to output dimension collapse.
To formally characterize this collapse, we derive the best prediction attainable within the collapsed subspace. This best prediction establishes a lower bound for prediction error under output dimension collapse. Let $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$ denote the subspace defined by linear combinations of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. The vector $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\in\mathcal{A}$ that is closest to the true latent feature vector $\bm{y}_{\mathrm{te}}$ can be derived analytically as:
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ (6)
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
where $A^{+}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see A.2 for details). This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Figure 2B).
Crucially, the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
$\forall\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A},\quad\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}\leq\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}.$ (8)
This inequality shows that no prediction $\hat{\bm{y}}_{\mathrm{te}}$ can ever outperform the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$. Moreover, the prediction error $\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ can be decomposed into two components:
$\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}=\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}+\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}.$ (9)
The first term $\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}$ corresponds to the within-subspace error, and the second term $\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ corresponds to the out-of-subspace error. When $n<d_{\mathrm{out}}$, the out-of-subspace error is generally non-zero, implying an irreducible error dictated by the training outputs. This error originates from the training outputs being insufficient to cover the universe of possible images. We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and actual predictions from brain activity are close to this analytic limit.
4 Empirical study: ODC on real data
In the previous section, we showed that naive multivariate linear regression traps in output dimension collapse at small data scales. This section examines how this collapse affects brain-to-image reconstruction in practice. Specifically, we first analyze the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ under various data scales and then compare $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the actual prediction from brain $\hat{\bm{y}}_{\mathrm{te}}$.
Our experiments mainly follow the method of Shen et al. (2019), which employs a Translator–Generator framework to reconstruct images from brain activity. The original method adopts a sparse linear regression model as the translator, and the generator iteratively optimizes the reconstructed image via a CNN-based process. Here, we replace the sparse translator with naive multivariate linear regression to observe the effect of output dimension collapse. For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with minimal prior influence, following the protocol of Shen et al. (2019) for iterative image optimization. We also evaluated the method of Ozcelik and VanRullen (2023), which adopts naive multivariate linear regression as the translator (Figure 2). We selected the Deeprecon dataset (Shen et al., 2019), designed explicitly for brain-to-image reconstruction. The training set comprises 1,200 natural images from ImageNet (Deng et al., 2009), and the test set consists of 50 natural images from different categories (unseen during training) and 40 artificial shapes drawn entirely from outside the training domain. Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling a rigorous zero-shot evaluation.
To see how the data scale drives output dimension collapse, we progressively reduce the training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition. Figure 3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$. As $n$ decreases, the error grows, indicating that the dimension of the training subspace is shrinking and cannot fully capture the variability in the true latent features. Figure 3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale. While the reconstruction from the uncollapsed true feature $\mathcal{G}(\bm{y}_{\mathrm{te}})$ is nearly perfect, the quality of the reconstruction from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ worsens noticeably as $n$ decreases. These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through output dimension collapse.
Figure 3: The best predictions for different training data sizes. (A) The best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across training data sizes 1200, 600, 300, and 150. The vertical axis represents the error of the best prediction, and the horizontal axis represents the sample index sorted by the error at n=n=150. (B) Representative reconstructions from the best predictions for different training data sizes. From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best prediction values $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples. The left four columns show natural images, and the right two columns show artificial shape images.
To assess how output dimension collapse limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$. Figure 4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$. For natural images, the best prediction error accounts for 70% of the total prediction error, and even for artificial images, the proportion reaches 40% (left panel). Examining individual samples reveals a strong correlation between the best and brain prediction errors (right panel). Samples that lie farther from the subspace (i.e., with larger best prediction errors) exhibit correspondingly larger total prediction errors. These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of output dimension collapse. Figure 1 presents the corresponding results with subsampled training sets. Figure 4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Both reconstructions are similarly blurred, markedly different from the “true” reconstruction. We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Figure 2). As a reference point, Shen et al.’s original method makes different design choices (e.g., a sparse translator) and can yield sharper images.
Figure 4: Comparison of the best prediction and the brain prediction. (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. The bottom row shows the latest method reconstructions (Shen et al., 2019). The left four columns show natural images, and the right two columns show artificial shape images.
These results confirm that naive multivariate linear regression suffers severely from output dimension collapse due to small data scales. Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction. In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions. These results underscore that naive regression is incompatible with the fundamental structure of reconstruction data: the vast image space and the limited training data.
5 Leveraging sparse brain-to-feature structure
In the previous sections, we demonstrated that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli. This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data. In contrast, brain-to-feature mapping is known to be sparse, and several studies have reported improved prediction performance by exploiting this sparsity. In this section, we demonstrate that sparse linear regression, which leverages this structure, can yield low prediction errors even at small data scales. We first show that sparse regression is not subject to output dimension collapse. We then formalize a data generation process and estimate the prediction error characterized by a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
In brain-to-image reconstruction, employing “brain-like” latent features renders the brain-to-feature mapping sparse. Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000). Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Figure 5A).
Figure 5: Sparse brain-to-feature mapping. (A) If the brain and the latent features exhibit local or selective activity, their relationship should become sparse mapping. (B) Data generation process that reflects the sparse structure of the brain-to-feature mapping. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Sparse regression models can effectively exploit the sparse brain-to-feature mapping. The model selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters. Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)). Importantly, these models are not subject to output dimension collapse at small data scales ($n<d_{\mathrm{out}}$). Naive regression uses a single shared set of input variables to predict all output dimensions, which implies $\mathrm{rank}(\hat{W})\leq n$. In contrast, a sparse model can select a different subset of input variables for each output dimension, allowing $\mathrm{rank}(\hat{W})>n$. Conceptually, ODC results from a mismatch between the implicit parameter reduction of the naive model and the data structure. When the true mapping is sparse, explicit reduction through variable selection that matches this structure should be preferred. Theoretically, sparse regression can achieve accurate prediction with small data scales by leveraging sparsity (Wainwright, 2009a; Donoho, 2006; Fan and Lv, 2008). Empirically, reconstruction methods that employ brain-like latent features and sparse regression models have achieved accurate zero-shot reconstruction even with limited training data (Miyawaki et al., 2008; Shen et al., 2019).
To engage in a deeper discussion, we formalize a data generation process that reflects the sparse structure of the brain-to-feature mapping (Figure 5B). Let us assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$ for simplicity. Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated as follows:
$\bm{y}^{(i)}$ $=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)},$ (10)
where the brain activity $\bm{x}^{(i)}$ is generated by sampling $\bm{x}^{(i)}\sim\mathcal{N}(\bm{0},I_{d})$, the noise $\bm{\epsilon}^{(i)}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{d})$, and the teacher weight matrix $W\in\mathbb{R}^{d\times d}$ is column-wise $s$-sparse:
$W_{j,k}=\begin{cases}\frac{1}{\sqrt{s}}&j\in S_{k},\\ 0&j\notin S_{k},\end{cases}\quad(j=1,2,\ldots,d),$ (11)
with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices. We denote the non-zero weight ratio by $a=s/d$, reflecting the fraction of input dimensions that contribute to each output dimension. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Under this generative setting (Figure 5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support. Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected. The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
$R\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\},$ (12)
with
$\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},\qquad\tau^{2}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\qquad\rho=\frac{p_{\mathrm{sel}}}{\gamma}$ (13)
(see A.3 for the derivation). Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$. In case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection. This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods. Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013). Figure 6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$. With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases. Along the diagonal $p_{\mathrm{sel}}=\pi$ (random selection), the error does not decrease, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
Figure 6: Prediction error of sparse regression. Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right). The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$. The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection. The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
Sparse regression reduces prediction error by leveraging the sparse brain-to-feature mapping, a structure that arises when brain-like latent features are employed. By formalizing this sparse structure, we have shown that appropriate variable selection can lead to accurate prediction even at small data scales.
6 Quantifying hit rate and prediction error in practical sparse regression
In the previous section, we analyzed how the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ jointly determine the prediction error in sparse regression. However, the selection algorithm and the data determine these two quantities. In this section, we focus on a concrete variable selection procedure and derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Furthermore, by comparing the prediction errors by our theory with those from simulations, we confirm the accuracy of the analysis and its applicability.
Among the various variable selection methods, we focus on the simple correlation-based (akin to SIS (Fan and Lv, 2008)) as a concrete method (Figure 7). This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data. For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $p_{\mathrm{sel}}$ fraction of input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$. A separate ridge regression uses only the selected inputs to predict each output dimension. Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
Figure 7: Correlation-based filter. The depth labeled $n$ is the sample axis: each slice represents one training sample. For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}$ by absolute correlation, and fit the model using only the selected inputs.
Under the data generation process introduced in Section 5 (Figure 5B), we can derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. We obtain the following relation for large $d$:
$\pi\approx\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right),$ (14)
with
$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\qquad\alpha=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (15)
where $\Phi$ is the cumulative distribution function of the standard normal distribution. A detailed derivation is provided in A.4. This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving (14) for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Figure 8 visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$. For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$. In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
Figure 8: Hit rate versus selection rate with a correlation-based filter. Each curve shows the estimated prediction error $R$ for different non-zero weight ratios $a$. The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$. Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter. Under the assumed data generation process (Figure 5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using (12) and (14). We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions. For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under (12) and (14). Figure 9 summarizes the theoretical estimates. For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data. For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller. These results demonstrate that a practical sparse regression model can translate the sparse structure of the data into substantial accuracy gains and that, when the mapping is sufficiently sparse, it can achieve accurate zero-shot prediction even at small data scales.
Figure 9: The estimated prediction error from theory. The black line shows the prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed. The colored lines show the prediction error of sparse regression for different non-zero weight ratios $a$. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$.
To assess the accuracy and applicability of our theory, we conduct simulations under four data generation processes (Figure 10). Specifically, we consider: (A) a baseline that exactly matches the assumptions used in our analysis (Figure 5B); (B) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (C) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect correlation structure in neural activity; and (D) input noise representing neural variability and measurement noise (see A.5 for details). Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors. For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales. This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation. These results indicate that our framework provides reliable and informative guidance in practical settings.
Figure 10: The simulated prediction error under four data-generating processes. Each plot shows the simulated prediction error $R$ of sparse regression under four different data-generating processes: (A) Baseline (the same as the setting in our theory), (B) Gaussian-distributed non-zero weights, (C) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (D) Input noise. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$. Each colored line shows the prediction error for a different non-zero weight ratio $a$.
7 Discussion
In this study, we theoretically analyzed how translator models behave under small data scales. We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluate its impact on the reconstruction pipeline that employs a naive translator. We then mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters. Our results show that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure. These findings imply that sparse regression does not suffer from output dimension collapse; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator. Although this best prediction minimizes latent-feature error, it does not always yield the most accurate reconstruction. Even so, it is evident that the output dimension collapse caused by small data scales prevents the model from accurately capturing the true latent features, thereby compromising the quality of the reconstructed images. Moreover, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al. (2025). The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment. Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
In Section 5 and Section 6, we evaluated sparse linear regression under the idealized assumptions that model real-world data. Although real datasets do exhibit a performance boost from sparse regression, the improvement is typically more modest than in our simulations. We performed additional simulations under a range of more realistic conditions, yet some discrepancies with actual data are likely to remain. First, our simulations cover relatively high SNR regimes. Today’s brain data contains more noise (Allen et al., 2022; Hebart et al., 2023), suggesting that the effective noise level is higher than the values used in our simulations. Although our theory is applicable regardless of the noise level, obtaining a small prediction error requires sufficiently low noise. Second, the latent features used in practice may not fully satisfy the sparsity assumption. While our simulations considered several sparsity levels, it is challenging to estimate the degree of sparsity in real data. Such sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in practice, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely meet this ideal. Improvements in neural recording denoising techniques and the design of AI feature spaces that empirically approximate the brain’s representations will likely enable zero-shot reconstruction at small data scales.
Although we examined only linear translators, recent reconstruction methods sometimes deploy nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023, 2024). For instance, Chen et al. (2023) employed Masked Brain Modeling, embedding structural priors of brain activity into the translator in an attempt to improve prediction accuracy. Although nonlinear networks can encode such inductive biases, output dimension collapse remains a concern. In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid output dimension collapse. Empirically, Shirakawa et al. (2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena. A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction. In the latest method, as described in Section 3, the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses. More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024). These models enable us to generate photorealistic images even from inaccurate latent features. However, this approach carries the risk of introducing information not actually present in the brain’s representation. In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity. The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language. For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020). When solving regression at small data scales, output dimension collapse is a critical concern. Sparse regression is effective in avoiding collapse; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021). Although we do not treat this situation here, determining how such selection concentration shapes prediction remains an important direction for future work.
7.1 Conclusion
The significance of our study lies in its dual contribution: First, we show that naive linear regression suffers from output dimension collapse at small data scales, restricting generalization beyond the training data. Second, we demonstrate that sparse linear regression can achieve accurate prediction at small data scales by mathematically characterizing the prediction error. These findings provide a practical guideline for advancing image reconstruction without relying on large-scale data.
8 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (GPT-4o, o3, 5) in order to improve readability and English translation. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
9 Acknowledgements
We thank our laboratory team, especially Ken Shirakawa and Hideki Izumi, for their invaluable feedback and insightful suggestions on the manuscript. This work was supported by Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705, JP25H00450 to Y.K., and JP21K17821 to Y.N.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, JPMJCR22P3 to Y.K., and BOOST grant JPMJBS2407 to K.O.), New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.), and Guardian Robot Project, RIKEN.
Appendix A Supplementary Materials
A.1 Output subspace restriction in kernel linear regression and PLS
This section shows that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
Kernel linear regression.
Let $k(\cdot,\cdot)$ be a positive semidefinite kernel and $K\in\mathbb{R}^{n\times n}$ the Gram matrix with entries $K_{ij}=k(\bm{x}^{(i)},\bm{x}^{(j)})$. For a test input $\bm{x}_{\mathrm{te}}$, define $\bm{k}_{\mathrm{te}}=[k(\bm{x}_{\mathrm{te}},\bm{x}^{(1)}),\ldots,k(\bm{x}_{\mathrm{te}},\bm{x}^{(n)})]^{\top}$. Kernel linear regression fits functions in the RKHS and, by the representer theorem, the multi-output predictor takes the matrix form:
$\hat{\bm{y}}_{\mathrm{te}}\;=\;A^{\top}\bm{k}_{\mathrm{te}}\quad\text{with}\quad A\;=\;(K+\lambda I)^{-1}Y_{\mathrm{tr}}.$
Therefore
$\hat{\bm{y}}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}\bm{m}=\sum_{i=1}^{n}m_{i}\,\bm{y}^{(i)},$
where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$. Thus, every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
Partial least squares (PLS).
We consider multivariate PLS with $r$ components on $(X_{\mathrm{tr}},Y_{\mathrm{tr}})$. Two standard variants are NIPALS and SIMPLS, which differ in how they construct the training score matrix $T\in\mathbb{R}^{n\times r}$ and the out-of-sample score map $t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{r}$, yet the prediction formula below does not depend on which algorithm produced $T$ and $t(\cdot)$.
Given $T$, the Y-side loading follows from least squares:
$Q^{\top}\;=\;\underset{B\in\mathbb{R}^{r\times d_{\mathrm{out}}}}{\operatorname{argmin}}\|\,Y_{\mathrm{tr}}-TB\,\|_{F}^{2}\;=\;(T^{\top}T)^{-1}T^{\top}Y_{\mathrm{tr}}.$
For a test input $\bm{x}_{\mathrm{te}}$, PLS predicts
$\hat{\bm{y}}_{\mathrm{te}}\;=\;t(\bm{x}_{\mathrm{te}})^{\top}Q^{\top}\;=\;t(\bm{x}_{\mathrm{te}})^{\top}(T^{\top}T)^{-1}T^{\top}Y_{\mathrm{tr}}\;=\;Y_{\mathrm{tr}}^{\top}\,\bm{m},$
where $\bm{m}\;=\;T(T^{\top}T)^{-1}t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{n}$. Thus every prediction $\hat{\bm{y}}_{\mathrm{te}}$ is a linear combination of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$. This conclusion holds for NIPALS and SIMPLS alike, regardless of the specific construction of $T$ and $t(\cdot)$.
A.2 The best prediction
The best prediction in naive multivariate linear regression is derived analytically as follows:
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
where $A^{+}$ is the Moore-Penrose pseudo-inverse of $A$.
Proof.
Since $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$, we can express the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ as
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$
$=\underset{\bm{m}\in\mathbb{R}^{n}}{\operatorname{argmin}}\lVert Y_{\mathrm{tr}}^{\top}\bm{m}-\bm{y}_{\mathrm{te}}\rVert^{2}$
The normal equation obtains the solution:
$Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\bm{m}=Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
Thus, the best prediction is
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
A.3 Prediction error for sparse linear regression
We quantify the prediction error when the sparse regression retains a fraction $p_{\mathrm{sel}}$ of the inputs and captures a fraction $\pi$ of the truly relevant ones. We focus on the asymptotic regime where both the sample size $n$ and the dimensionality $d$ are taken to infinity, while the data scale $n/d$ stays finite. By applying the risk formula for ridge regression by Dicker (2016) to the post-selection problem, we quantify the prediction error of sparse regression.
Consider the linear model that remains after variable selection for a single coordinate $y_{k}$ of $\bm{y}$. Split every input vector $\bm{x}^{(i)}\in\mathbb{R}^{d}$ into two parts: the selected part $\bm{\xi}^{(i)}_{k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and the unselected part $\bm{\zeta}^{(i)}_{k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$, and similarly the weight vector $\bm{w}_{k}=W_{:,k}\in\mathbb{R}^{d}$ into the corresponding $\bm{u}_{k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and $\bm{v}_{k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$. With these notations
$y_{k}^{(i)}$ $=\bm{w}_{k}^{\top}\bm{x}^{(i)}+\epsilon_{k}^{(i)}$ (16)
$=\bm{u}_{k}^{\top}\bm{\xi}^{(i)}_{k}+\bm{v}_{k}^{\top}\bm{\zeta}^{(i)}_{k}+\epsilon_{k}^{(i)}$ (17)
$=\bm{u}_{k}^{\top}\bm{\xi}^{(i)}_{k}+\eta_{k}^{(i)}$ (18)
where $\eta_{k}^{(i)}=\bm{v}_{k}^{\top}\bm{\zeta}^{(i)}_{k}+\epsilon_{k}^{(i)}$ is the combined noise term. The selected part of weight vector $\bm{u}_{k}$ contains $\pi s$ non-zero entries, and the unselected part $\bm{v}_{k}$ contains $(1-\pi)s$ non-zero entries, hence $\|\bm{u}_{k}\|^{2}=\pi$ and $\|\bm{v}_{k}\|^{2}=1-\pi$. Further, since $\bm{x}^{(i)}$ is generated from $\mathcal{N}(\bm{0},I_{d})$ and independent of $\bm{\epsilon}^{(i)}$, the combined noise satisfies
${\eta}_{k}^{(i)}\sim\mathcal{N}(0,\sigma_{\mathrm{eff}}^{2}),$ (19)
with $\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2}$, and it is independent of $\bm{\xi}^{(i)}_{k}$. The post-selection problem is therefore a standard linear regression with effective dimension $d_{\mathrm{sel}}$ and effective noise variance $\sigma_{\mathrm{eff}}^{2}$.
The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: in the limit $n,d_{\mathrm{sel}}\to\infty$ with $n/d_{\mathrm{sel}}<\infty$, the risk of the ridge estimator with the optimal regularization parameter is
$\mathbb{E}\|\hat{\bm{u}}_{k}-\bm{u}_{k}\|^{2}\approx\frac{\sigma_{\mathrm{eff}}^{2}}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr],$ (20)
with
$\tau^{2}=\frac{\|\bm{u}_{k}\|^{2}}{\sigma_{\mathrm{eff}}^{2}}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\quad\rho=\frac{d_{\mathrm{sel}}}{n}=\frac{p_{\mathrm{sel}}}{\gamma},\quad\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},$ (21)
where $\gamma=n/d$ is the data scale. For the test data $(\bm{x}_{\mathrm{te}},\bm{y}_{\mathrm{te}})$, we can split $\bm{x}_{\mathrm{te}}$ into $\bm{\xi}_{\mathrm{te},k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and $\bm{\zeta}_{\mathrm{te},k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$ in the same way as above. The $k$-th output coordinate of the test data $\bm{y}_{\mathrm{te}}$ is similarly expressed as
$y_{\mathrm{te},k}=\bm{u}_{k}^{\top}\bm{\xi}_{\mathrm{te},k}+\eta_{\mathrm{te},k},$ (22)
where $\eta_{\mathrm{te},k}\sim\mathcal{N}(0,\sigma_{\mathrm{eff}}^{2})$ is independent of $\bm{\xi}_{\mathrm{te},k}$. The learned weight on the unselected part $\bm{\zeta}_{\mathrm{te},k}$ is zero, so the prediction for the $k$-th output coordinate is
$\hat{y}_{\mathrm{te},k}$ $=\hat{\bm{w}}_{k}^{\top}\bm{x}_{\mathrm{te}}$ (23)
$=\hat{\bm{u}}_{k}^{\top}\bm{\xi}_{\mathrm{te},k},$ (24)
where $\hat{\bm{u}}_{k}$ is the learned weight on the selected part $\bm{\xi}_{\mathrm{te},k}$. Therefore, the prediction error $R$ is
$R$ $=\frac{1}{d}\mathbb{E}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ (25)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|\hat{y}_{\mathrm{te},k}-y_{\mathrm{te},k}\|^{2}\right]$ (26)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|(\hat{\bm{u}}_{k}-\bm{u}_{k})^{\top}\bm{\xi}_{\mathrm{te}}-\eta_{\mathrm{te},k}\|^{2}\right]$ (27)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|(\hat{\bm{u}}_{k}-\bm{u}_{k})^{\top}\bm{\xi}_{\mathrm{te}}\|^{2}+\|\eta_{\mathrm{te},k}\|^{2}\right]$ (28)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|\hat{\bm{u}}_{k}-\bm{u}_{k}\|^{2}+\sigma_{\mathrm{eff}}^{2}\right]$ (29)
$\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\}.$ (30)
This expression gives the prediction error of sparse regression with a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
A.4 Selection rate-hit rate relation for correlation-based filter
We relate the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ of the correlation-based filter to characterize the performance of the practical filter. We work in the same asymptotic regime as in the preceding analysis: $n,d\to\infty$ with $n/d<\infty$. All $\approx$’ relations are asymptotic equalities in this limit.
In the following, we express the column slices of data matrices $\left(X_{\mathrm{tr}}\right)_{:,j},\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ as $\mathbf{x}_{j}$ or $\mathbf{y}_{k}$. We intentionally use different symbol from $\bm{x}^{(i)}$ or $\bm{y}^{(i)}$ which indicates $i$-th single instances. For each output coordinate $k\in\{1,\ldots,d\}$, we can rewrite the data generation process as
$\mathbf{y}_{k}$ $=X_{\mathrm{tr}}W_{:,k}+\mathbf{\epsilon}_{k}$ (31)
$=\sum_{j=1}^{d}W_{jk}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (32)
$=\frac{1}{\sqrt{s}}\sum_{j\in\mathcal{S}_{k}}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (33)
where $W_{:,k}$ is the $k$-th column of the teacher weight matrix $W$, $\mathbf{\epsilon}_{k}\in\mathbb{R}$ is the $k$-th output noise vector, and $\mathcal{S}_{k}=\left\{j\mid W_{jk}\neq 0\right\}$ is the informative index set for the $k$-th output coordinate. Within this data generation process, $\mathbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\mathbf{\epsilon}_{k}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{n})$ hold.
The correlation-based filter ranks each input variable $j$ according to its absolute correlation with the output $k$:
$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$ (34)
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Because $\textbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\textbf{y}_{k}\sim\mathcal{N}(\bm{0},(1+\sigma^{2})I_{n})$ hold, the means and standard deviations satisfy $\hat{\mu}_{\mathrm{x}_{j}}\approx 0,\hat{\sigma}_{\mathrm{x}_{j}}\approx 1,\hat{\mu}_{\mathrm{y}_{k}}\approx 0,\hat{\sigma}_{\mathrm{y}_{k}}\approx\sqrt{1+\sigma^{2}}$. Thus, the correlation coefficient simplifies to
$r_{jk}\approx\frac{\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle}{n\sqrt{1+\sigma^{2}}}=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle\times\mathrm{const}.$ (35)
Hence, the ranking by the absolute correlation coefficient $|r_{jk}|$ is almost equivalent to the ranking by the absolute inner product $|z_{jk}|=\left|\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle\right|$.
We first derive the expectation and variance of the inner product $z_{jk}=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle$ for informative inputs ($j\in\mathcal{S}_{k}$) and uninformative inputs ($j\notin\mathcal{S}_{k}$). For an informative input $j\in\mathcal{S}_{k}$, we can write
$z_{jk}$ $=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle$ (36)
$=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (37)
$=\frac{1}{\sqrt{s}}\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (38)
Independence of the Gaussian components yields
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (39)
$=\frac{n}{\sqrt{s}}+0+0=\frac{n}{\sqrt{s}},$ (40)
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{s}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (41)
$=\frac{2n}{s}+\frac{(s-1)n}{s}+n\sigma^{2}=n(\frac{s+1}{s}+\sigma^{2}).$ (42)
Similarly, for an uninformative input $j\notin\mathcal{S}_{k}$, we find
$z_{jk}$ $=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (43)
$=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (44)
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (45)
$=0+0=0,$ (46)
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\sum_{t\in\mathcal{S}_{k}}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (47)
$=n(1+\sigma^{2}).$ (48)
Next, we derive the relation between the hit rate $\pi$ and the selection rate $p_{\mathrm{sel}}$ of the correlation-based filter. Fix an output $k$ and let the true support be $\mathcal{S}_{k}=\{j\mid w_{jk}\neq 0\}$ with $|\mathcal{S}_{k}|=s=ad$. For a common threshold $t_{k}>0$, define the selection set $\hat{\mathcal{S}}_{k}(t_{k})=\{j\mid|z_{jk}|>t_{k}\}$. Recall the selection rate $p_{\mathrm{sel}}=\frac{|\hat{\mathcal{S}}_{k}(t_{k})|}{d}=\Pr(j\in\hat{\mathcal{S}}_{k}(t_{k}))$ and the hit rate $\pi=\Pr(j\in\hat{\mathcal{S}}_{k}(t)\mid j\in\mathcal{S}_{k})$. Define the null retention probability $\pi_{-}:=\Pr(j\in\hat{\mathcal{S}}_{k}(t_{k})\mid j\notin\mathcal{S}_{k})$. Noting that $a=\Pr(j\in\mathcal{S}_{k})$, we obtain the identity
$p_{\mathrm{sel}}$ $=a\,\pi+(1-a)\,\pi_{-}.$ (49)
We proceed by approximating $\pi$ and $\pi_{-}$ and then eliminating the threshold. Denote by $\Phi(\cdot)$ and $\Phi^{-1}(\cdot)$ the CDF and inverse-CDF of the standard normal distribution. Writing $u_{k}:=t_{k}/\sqrt{n}$ and $\Delta:=\sqrt{n/s}=\sqrt{\gamma/a}$ with data scale $\gamma=n/d$, we obtain
$\pi$ $=\Pr(|z_{jk}|>t_{k}\mid j\in\mathcal{S}_{k})$ (50)
$=\Phi\!\left(-\frac{t_{k}-\frac{n}{\sqrt{s}}}{\sqrt{\,n\!\left(\frac{s+1}{s}+\sigma^{2}\right)}}\right)+\Phi\!\left(-\frac{t_{k}+\frac{n}{\sqrt{s}}}{\sqrt{\,n\!\left(\frac{s+1}{s}+\sigma^{2}\right)}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (51)
$=\Phi\!\left(-\frac{u_{k}-\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)+\Phi\!\left(-\frac{u_{k}+\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (52)
$=\Phi\!\left(-\frac{u_{k}-\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\Phi\!\left(-\frac{u_{k}+\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (53)
(51) follows from Berry-Esseen theorem, and (53) follows from
$\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)=\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(1/s\right).$ (54)
Similarly,
$\pi_{-}$ $=\Pr(|z_{jk}|>t_{k}\mid j\notin\mathcal{S}_{k})$ (55)
$=2\,\Phi\!\left(-\frac{t_{k}}{\sqrt{\,n(1+\sigma^{2})\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (56)
$=2\,\Phi\!\left(-\frac{u_{k}}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (57)
Combining (53) and (57) with (49) yields the two-equation relation
$p_{\mathrm{sel}}$ $=a\!\left[\Phi\!\left(-\tfrac{u_{k}-\Delta}{\sqrt{1+\sigma^{2}}}\right)+\Phi\!\left(-\tfrac{u_{k}+\Delta}{\sqrt{1+\sigma^{2}}}\right)\right]$
$\quad+(1-a)\,2\,\Phi\!\left(-\tfrac{u_{k}}{\sqrt{1+\sigma^{2}}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right),$ (58)
$\pi$ $=\Phi\!\left(-\tfrac{u_{k}-\Delta}{\sqrt{1+\sigma^{2}}}\right)+\Phi\!\left(-\tfrac{u_{k}+\Delta}{\sqrt{1+\sigma^{2}}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right),$ (59)
which links $p_{\mathrm{sel}}$ and $\pi$ through the common threshold $u_{k}>0$. Eliminating $u_{k}$ gives, for $a\in(0,1)$,
$\pi\approx\Phi\!\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\!\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right)$ (60)
where
$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\quad\alpha=\frac{\Delta}{\sqrt{1+\sigma^{2}}}=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (61)
while in the boundary case $a=1$ the relation reduces to $p_{\mathrm{sel}}=\pi$.
A.5 Detailed Simulation Settings
We conducted simulations under the following four generative conditions:
  • (A)
    Baseline: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. Each non-zero entry in the teacher matrix $W$ is set to $1/\sqrt{s}$, with the remaining entries set to 0. The input vector $\bm{x}[{(i)}]$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
  • (B)
    Gaussian non-zero weights: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as in (A), but the non-zero entries in the teacher matrix are sampled from a standard normal distribution $\mathcal{N}(0,\frac{1}{s})$. Inputs $\bm{x}^{(i)}$ are independently drawn from $\mathcal{N}(0,1)$, as in (A).
  • (C)
    Correlated signal: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as in (A), with the same binary non-zero structure in the teacher matrix $W$. The input vector $\bm{x}^{(i)}$ is sampled from a zero-mean multivariate normal distribution with a Toeplitz covariance matrix $\Sigma_{ij}=\rho^{|i-j|}$, where $\rho=0.5$.
  • (D)
    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)},\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. The teacher matrix $W$ has the same binary non-zero structure as in (A). The input vector $\bm{z}^{(i)}$ is sampled from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
The simulations are performed with the following parameters: the dimensionality $d=1000$ and the noise level $\sigma=0.1$.
Appendix B Supplementary Figures
Figure 1: Comparison of the best prediction and the brain prediction for subsampled training sets. Each panel shows the same analysis as in Figure 4 with subsampled training sets of 1200, 600, 300, and 150 samples. Each dot represents one image, and the dotted line indicates where the two errors are equal. Blue for natural images, orange for artificial shapes. As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
Figure 2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023). (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Similar results are observed: the brain prediction and best prediction are close to each other but far from the true feature.

### v2

@@
-Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales
+Overcoming Output Dimension Collapse: When Sparsity Enables Zero-shot Brain-to-image Reconstruction at Small Data Scales
 Abstract
 Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images.
-Achieving such reconstruction with limited training data requires generalization beyond the training set, a task known as zero-shot prediction.
-Despite its importance, we still lack theoretical guidelines for achieving efficient and accurate reconstruction.
-In this paper, we provide a theoretical analysis of two widely used models for translating brain activity to latent image features.
-We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize how each model behaves across data scales.
+A key challenge in this task is data scarcity: a translator that maps brain activity to latent image features is trained on a limited number of brain-image pairs, making the translator a bottleneck for zero-shot reconstruction beyond the training stimuli.
+In this paper, we provide a theoretical analysis of two translator designs widely used in recent reconstruction pipelines: naive multivariate linear regression and sparse multivariate linear regression.
+We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize the behavior of each model across data scales.
 We first show that the naive linear regression model, which uses a shared set of input variables for all outputs, suffers from “output dimension collapse” at small data scales, restricting generalization beyond the training data.
-We then mathematically characterize the prediction error of the sparse linear regression model by deriving formulas linking prediction error with data scale and other problem parameters.
-Leveraging the sparsity of the brain-to-feature mapping, this approach enables accurate zero-shot prediction even at small data scales without trapping in output dimension collapse.
-Our results provide a theoretical guideline for achieving zero-shot reconstruction and highlight the benefits of variable selection in brain-to-image reconstruction.
-keywords:
-Brain decoding , Brain-to-image reconstruction , Zero-shot prediction , Sparse linear regression
-[…]
-[…]
+We then analyze sparse linear regression models in a student–teacher framework and derive expressions for the prediction error in terms of data scale and other sparsity-related parameters.
+Our analysis clarifies when variable selection can reduce prediction error at small data scales by exploiting the sparsity of the brain-to-feature mapping.
+Our findings provide quantitative guidelines for diagnosing output dimension collapse and for designing effective translators and feature representations for zero-shot reconstruction.
 1 Introduction
 Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images.
-To uncover neural representations and move toward practical applications in medicine and industry, it will be necessary to reconstruct a broad range of subjective visual experiences.
-Nevertheless, a fundamental constraint governs this task: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training.
-To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, namely zero-shot prediction (Shirakawa et al., 2025).
+To uncover neural representations and move toward practical applications in medicine and industry, we need reconstruction methods that generalize to a broad range of subjective visual experiences.
+A key challenge in this task is data scarcity: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training.
+To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, i.e., zero-shot prediction (Shirakawa et al., 2025).
 Despite its importance, we still lack theoretical guidelines for achieving accurate reconstruction under these constraints.
-Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1).
-A translator module converts brain activity into latent features that represent the image content, and a generator module transforms these features into reconstructed images.
-This paper focuses on the translator module, trained on a limited pairs of brain activity and latent features.
-We quantify this limitation by the data scale, the ratio of the number of training pairs to the dimensionality of the latent features.
-To realize zero-shot reconstruction, the translator must be able to predict valid latent features for unseen stimuli at small data scales.
+Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Fig.
+1).
+In this framework, the translator converts brain activity into latent features that represent the visual content, and the generator transforms these features into reconstructed images.
+Latent features are typically high-dimensional to represent a broad spectrum of visual stimuli.
+Recent pipelines often leverage powerful generators, but the translator must be trained using only a limited number of brain-feature pairs.
+We quantify this data limitation using the data scale $n/d_{\mathrm{out}}$, where $n$ is the number of training pairs and $d_{\mathrm{out}}$ is the latent feature dimensionality.
+The translator becomes a key bottleneck for zero-shot reconstruction because it must predict valid latent features for unseen stimuli when $n/d_{\mathrm{out}}<1$.
+A common translator design is what we call naive multivariate linear regression, which uses a shared set of input variables to predict all output dimensions (e.g., ridge regression).
+This design is straightforward and low-cost, making it a common choice in recent studies (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023).
+Another approach is sparse multivariate linear regression, which performs variable selection for each output dimension (e.g., Lasso, ARD, filter methods).
+Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
+Theoretical results suggest that sparse regression can be sample-efficient when the underlying mapping is sparse (Donoho, 2006; Wainwright, 2009b).
+Notably, these linear translators are not only classical baselines but are still central components in many recent reconstruction pipelines.
 Figure 1: Translator–Generator pipeline.
@@
 A naive multivariate linear regression model and a sparse multivariate linear regression model are commonly used as translators.
-Whereas a naive model uses a single shared set of input variables to predict all output dimensions, a sparse model allows each output dimension to select its own subset of input variables.
-Two linear regression models are widely used as translators: naive multivariate linear regression and sparse multivariate linear regression.
-A naive multivariate linear regression model, exemplified by ridge regression, uses a single shared set of input variables to predict all output dimensions and has been adopted in several high-profile studies (Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023).
-This strategy is straightforward and low-cost, yet Shirakawa et al.
-(2025) reports that such models fail to generalize beyond the training domain.
-They attribute this failure to “output dimension collapse,” in which predictions are confined to the span of the training outputs under dataset bias.
-In contrast, a sparse multivariate linear regression model allows each output dimension to select its own subset of input variables.
-Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
+A naive model uses a shared set of input variables to predict all output dimensions, and a sparse model performs variable selection for each output dimension.
+In brain-to-image reconstruction, Shirakawa et al.
+(2025) pointed out a phenomenon in naive multivariate linear translators called “output dimension collapse” (ODC).
+They reported that several high-profile reconstruction methods fail to generalize beyond the training domain.
+They attributed this failure to ODC, in which predictions are confined to the span of the training outputs in latent feature space under dataset bias.
+ODC is a challenge that must be addressed to achieve zero-shot prediction, but we still lack insight into when it occurs and how it can be overcome.
 In this paper, we provide a theoretical analysis of translator prediction under small data scales.
-First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales cause output dimension collapse in naive multivariate linear regression.
+First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales can cause ODC in naive multivariate linear regression.
 Using the mathematically derived best prediction under this collapse, we evaluate its impact on reconstruction pipelines that employ naive translators.
-Second, we examine how sparsity can enable accurate prediction at small data scales by characterizing the prediction error of naive and sparse translators.
-Our contributions are as follows:
-  • 1.
-    We show that reconstruction pipelines using naive linear regression suffer from output dimension collapse at small data scales.
-  • 2.
-    We mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters.
-  • 3.
-    We demonstrate that sparse linear regression can escape output dimension collapse and achieve accurate prediction at small data scales by leveraging sparse brain-to-feature structure.
+Second, we examine when sparsity can enable accurate prediction at small data scales by mathematically characterizing the prediction error of naive and sparse translators.
+The main contributions of this work are threefold: (1) a theoretical demonstration that naive multivariate linear regression exhibits output dimension collapse at small data scales; (2) an analytic diagnostic of the practical impact of ODC via the best prediction under collapse; and (3) a characterization of when sparsity enables accurate prediction at small data scales through quantitative error predictions for naive and sparse translators.
 2 Preliminaries
-Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1).
-In this framework, a module called the translator converts brain activity into latent features, and a module called the generator transforms these features into reconstructed images.
+Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Fig.
+1).
+In this framework, a module called the translator converts brain activity into latent features, and the generator transforms these features into reconstructed images.
 The translator is typically a linear regression model trained on paired brain activity and latent feature data.
 The generator is typically a pre-trained or pre-defined model that generates images from the predicted features.
-Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features (e.g., CNN, CLIP).
+Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features, such as Convolutional Neural Networks (CNNs) and Contrastive Language-Image Pretraining (CLIP).
 In this study, we focus on the translator module, which maps brain activity $\bm{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ to latent features $\bm{y}\in\mathbb{R}^{d_{\mathrm{out}}}$.
@@
 These methods differ in how they regularize or project the data, but still rely on shared inputs for all output dimensions.
-In contrast, sparse multivariate linear regression allows each output value to select its subset of input variables.
+Another approach is sparse multivariate linear regression, which performs variable selection for each output dimension.
 According to the classic taxonomy (Guyon and Elisseeff, 2003), selection strategies fall into three groups: wrapper methods (Kohavi and John, 1997), which iteratively search feature subsets by maximizing predictive performance; embedded methods (e.g., Lasso (Tibshirani, 1996), Automatic Relevance Determination (ARD) (MacKay, 1992)), which embed sparsity directly in the training objective; and filter methods (e.g., CFS (Hall, 2000), SIS (Fan and Lv, 2008) ), which rank inputs independently of the learner using simple scores (e.g., correlation, mutual information).
-In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and in practice, the runtime cost scales as wrapper $\gg$ embedded $\gg$ filter.
+In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and wrapper methods are typically the most computationally expensive, embedded methods have intermediate cost, and filter methods are usually the least expensive.
 Previous theoretical work demonstrates that sparsity is particularly beneficial when the sample size is smaller than the dimensionality of the data.
+We call an input–output mapping $s$-sparse if, for each output coordinate, the corresponding weight vector has at most $s$ non-zero entries (i.e., the output depends on only $s$ of the $d_{\mathrm{in}}$ input variables).
 For a truly $s$-sparse input-output mapping, information-theoretic bounds show that an ideal estimator can attain near-oracle prediction error once $n\gtrsim s\log(d_{\mathrm{in}}/s)$ (Wainwright, 2009a).
 The Lasso is the most thoroughly analyzed among practical sparse methods, and theory confirms it reaches essentially the same rate and achieves accurate prediction with $n=\mathcal{O}(s\log(d_{\mathrm{in}}))$ (Donoho, 2006; Wainwright, 2009b).
-The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ is exponentially large in $n$ (Fan and Lv, 2008).
-3 Output dimension collapse by small data scale
-Although both sparse and naive (non-sparse) models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023).
+The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ grows exponentially in $n$ (Fan and Lv, 2008).
+3 Output dimension collapse by small data scales
+Although both naive and sparse models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023).
 This section focuses on the naive translator and shows a critical limitation at small data scales.
-In brain-to-image reconstruction, the translator must map brain activity $\bm{x}$ onto a high-dimensional image feature space $\bm{y}$ while relying on a limited training sample.
-If we use a naive multivariate linear regression under such scenarios, a phenomenon “output dimension collapse” (ODC) can occur, whereby the predictions become restricted to the subspace determined by the training outputs (Shirakawa et al., 2025).
-This issue is especially critical for zero-shot prediction when the training subspace is low-dimensional, because the model inevitably compresses unfamiliar features into that subspace.
+Output dimension collapse (ODC) refers to a phenomenon in which predictions of naive multivariate linear regression are confined to a low-dimensional subspace spanned by the training outputs in the latent feature space.
+This restriction is especially critical to zero-shot prediction when the training-output subspace is low-dimensional, because unfamiliar features are inevitably compressed into that subspace.
 To understand why this collapse occurs, let us examine the predicted features of ridge regression.
-First, we learn a weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ by minimizing the sum of squared errors plus a regularization term:
+In ridge regression, the fitted weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ admits the following closed-form solution:
 $\hat{W}$ $=\underset{W}{\operatorname{argmin}}\sum_{i=1}^{n}\lVert\bm{y}^{(i)}-W^{\top}\bm{x}^{(i)}\rVert^{2}+\lambda\lVert W\rVert^{2}_{F}$ (1)
@@
 Clearly, the prediction $\hat{\bm{y}}_{\mathrm{te}}$ lies in the subspace spanned by the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$.
-In other words, no matter what the test input is, the predicted output is constrained to that subspace (Figure 2A).
+In other words, no matter what the test input is, the predicted output is constrained to that subspace (Fig.
+2A).
 Importantly, this property is not the familiar one that confines predictions to the span of the training inputs $\left\{\bm{x}^{(i)}\right\}_{i=1}^{n}$, but instead that it confines predictions to the span of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$.
 This subspace is determined solely by the training outputs and does not depend on the input brain data; given the same training outputs, the same property holds regardless of measurement modality (e.g., fMRI, MEG), subject, or decoding task (e.g., perception, imagery, or illusion).
-Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see A.1 for details).
+Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see Appendix B.1 for details).
 Figure 2: Output dimension collapse.
@@
 (B) The best prediction within the training feature subspace provides the lower bound for prediction error in naive multivariate linear regression.
-A key condition that leads to output dimension collapse is when the training outputs lie in a low-dimensional subspace.
+A key condition that leads to ODC is when the training outputs lie in a low-dimensional subspace.
 Shirakawa et al.
@@
 Even in the absence of explicit cluster structure, having $n<d_{\mathrm{out}}$ inherently forces all predictions into an at-most $n$-dimensional subspace.
-As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe output dimension collapse.
+As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe ODC.
 Consequently, regardless of whether the training outputs form visible clusters, small data scales alone can lead to the same fundamental difficulty in zero-shot prediction.
-From another perspective, output dimension collapse due to small data scale is an inevitable consequence of reducing the number of free parameters in the model.
+From another perspective, ODC by small data scales is an inevitable consequence of reducing the number of free parameters in the model.
 The model must learn a weight matrix with many parameters to predict high-dimensional latent features, but the available training samples are insufficient to identify all of them.
@@
 Naive linear regression accomplishes this by confining all predictions to the linear span of the training outputs.
-This implicit constraint is precisely what leads to output dimension collapse.
+This implicit constraint is precisely what leads to ODC.
 To formally characterize this collapse, we derive the best prediction attainable within the collapsed subspace.
-This best prediction establishes a lower bound for prediction error under output dimension collapse.
+This best prediction establishes a lower bound for prediction error under ODC.
 Let $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$ denote the subspace defined by linear combinations of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$.
@@
 $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ (6)
-$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
-where $A^{+}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see A.2 for details).
-This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Figure 2B).
-Crucially, the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
+$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
+where $A^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see Appendix B.2 for details).
+This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Fig.
+2B).
+Because it represents the best achievable prediction within this subspace, computing it in a given reconstruction pipeline allows us to assess how strongly ODC constrains its predictions directly.
+The best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
 $\forall\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A},\quad\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}\leq\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}.$ (8)
@@
 This error originates from the training outputs being insufficient to cover the universe of possible images.
-We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and actual predictions from brain activity are close to this analytic limit.
+We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and most of the observed prediction error can be explained by this irreducible out-of-subspace component.
 4 Empirical study: ODC on real data
-In the previous section, we showed that naive multivariate linear regression traps in output dimension collapse at small data scales.
-This section examines how this collapse affects brain-to-image reconstruction in practice.
+We showed that naive multivariate linear regression is prone to output dimension collapse at small data scales.
+We next examine how this collapse affects brain-to-image reconstruction in practice.
 Specifically, we first analyze the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ under various data scales and then compare $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the actual prediction from brain $\hat{\bm{y}}_{\mathrm{te}}$.
 Our experiments mainly follow the method of Shen et al.
-(2019), which employs a Translator–Generator framework to reconstruct images from brain activity.
-The original method adopts a sparse linear regression model as the translator, and the generator iteratively optimizes the reconstructed image via a CNN-based process.
-Here, we replace the sparse translator with naive multivariate linear regression to observe the effect of output dimension collapse.
-For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with minimal prior influence, following the protocol of Shen et al.
+(2019), which employs a Translator–Generator framework to reconstruct images from brain activity (see also Appendix B.3).
+To observe the effect of ODC, we replace their sparse translator with a naive multivariate linear regression model.
+For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with weak prior influence, following the protocol of Shen et al.
 (2019) for iterative image optimization.
-We also evaluated the method of Ozcelik and VanRullen (2023), which adopts naive multivariate linear regression as the translator (Figure 2).
+We also evaluated the method of Ozcelik and VanRullen (Ozcelik and VanRullen, 2023) as an additional baseline, which adopts naive multivariate linear regression as the translator.
 We selected the Deeprecon dataset (Shen et al., 2019), designed explicitly for brain-to-image reconstruction.
 The training set comprises 1,200 natural images from ImageNet (Deng et al., 2009), and the test set consists of 50 natural images from different categories (unseen during training) and 40 artificial shapes drawn entirely from outside the training domain.
-Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling a rigorous zero-shot evaluation.
-To see how the data scale drives output dimension collapse, we progressively reduce the training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition.
-Figure 3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$.
+Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling rigorous zero-shot evaluation.
+To see how the data scale drives ODC, we progressively reduce the number of training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition.
+Fig.
+3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$.
 As $n$ decreases, the error grows, indicating that the dimension of the training subspace is shrinking and cannot fully capture the variability in the true latent features.
-Figure 3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale.
+Fig.
+3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale.
 While the reconstruction from the uncollapsed true feature $\mathcal{G}(\bm{y}_{\mathrm{te}})$ is nearly perfect, the quality of the reconstruction from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ worsens noticeably as $n$ decreases.
-These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through output dimension collapse.
+These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through ODC.
 Figure 3: The best predictions for different training data sizes.
@@
 (B) Representative reconstructions from the best predictions for different training data sizes.
-From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best prediction values $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples.
+From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples.
 The left four columns show natural images, and the right two columns show artificial shape images.
-To assess how output dimension collapse limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$.
-Figure 4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$.
-For natural images, the best prediction error accounts for 70% of the total prediction error, and even for artificial images, the proportion reaches 40% (left panel).
+Figure 4: Comparison of the best prediction and the brain prediction.
+(A) Left: Percentage of the best prediction error relative to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$.
+Blue for natural images, orange for artificial shapes.
+The error bars denote the standard deviation across samples.
+Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$.
+Each dot represents one image, and the dotted line indicates where the two errors are equal.
+(B) Representative reconstructions generated from the best predictions and brain predictions.
+From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$.
+The bottom row shows reconstructions from the original method of Shen et al.
+(2019).
+The left four columns show natural images, and the right two columns show artificial shape images.
+To assess how ODC limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$.
+Fig.
+4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$.
+For natural images, the best prediction error accounts for 70% of the brain prediction error, and even for artificial images, the proportion reaches 40% (left panel).
 Examining individual samples reveals a strong correlation between the best and brain prediction errors (right panel).
 Samples that lie farther from the subspace (i.e., with larger best prediction errors) exhibit correspondingly larger total prediction errors.
-These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of output dimension collapse.
-Figure 1 presents the corresponding results with subsampled training sets.
-Figure 4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$.
+These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of ODC.
+Fig.
+A1 presents the corresponding results with subsampled training sets.
+Fig.
+4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$.
 Both reconstructions are similarly blurred, markedly different from the “true” reconstruction.
-We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Figure 2).
-As a reference point, Shen et al.’s original method makes different design choices (e.g., a sparse translator) and can yield sharper images.
-Figure 4: Comparison of the best prediction and the brain prediction.
+We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Fig.
+A2).
+As a reference point, Shen et al.’s original method (Shen et al., 2019) makes different design choices (e.g., a sparse translator) and can yield sharper images.
+These results confirm that naive multivariate linear regression suffers severely from ODC due to small data scales.
+Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction.
+In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions.
+These results underscore that naive regression is poorly suited to the fundamental structure of reconstruction datasets, in which a vast stimulus space must be learned from only a limited number of paired brain–image samples.
+5 Leveraging sparse brain-to-feature structure
+We have shown that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli.
+This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data.
+In contrast, the brain-to-feature mapping is commonly assumed to be sparse, and several studies have reported successful prediction performance under such sparsity assumptions (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
+We therefore theoretically analyze sparse regression to clarify how exploiting sparsity affects prediction.
+We first show that sparse regression is not forced into ODC in the same way as naive regression.
+We then analyze prediction error in a student–teacher setting for a broad class of sparse regression procedures.
+In brain-to-image reconstruction, using “brain-like” latent features makes it reasonable to assume a sparse brain-to-feature mapping.
+Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000).
+Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
+Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Fig.
+5A).
+Sparse regression provides an inductive bias that is well aligned with the brain-to-feature mapping.
+Sparse regression selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters.
+Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)).
+Figure 5: Sparse brain-to-feature mapping.
+(A) If the brain and the latent features exhibit local or selective activity, their mapping should become sparse.
+(B) Teacher model in our student–teacher framework, reflecting the sparse structure of the brain-to-feature mapping.
+The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
+[…]
+To clarify how the translator behaves under sparsity, we introduce a student–teacher model(Fig.
+5B).
+For simplicity, we assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$.
+Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated by a linear teacher as
+$\bm{y}^{(i)}$ $=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)},$ (10)
+where the brain activity $\bm{x}^{(i)}$ is generated by sampling $\bm{x}^{(i)}\sim\mathcal{N}(\bm{0},I_{d})$, the noise $\bm{\epsilon}^{(i)}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{d})$, and the teacher weight matrix $W\in\mathbb{R}^{d\times d}$ is column-wise $s$-sparse:
+$W_{j,k}=\begin{cases}\frac{1}{\sqrt{s}}&j\in S_{k},\\ 0&j\notin S_{k},\end{cases}\quad(j=1,2,\ldots,d),$ (11)
+with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices.
+We denote the non-zero weight ratio by $a=s/d$, which represents the fraction of input dimensions that contribute to each output dimension; the smaller $a$ is, the sparser the input–output mapping.
+Under this student–teacher setting (Fig.
+5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$.
+The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support.
+Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected.
+The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
+$R\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\},$ (12)
+with
+$\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},\qquad\tau^{2}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\qquad\rho=\frac{p_{\mathrm{sel}}}{\gamma}$ (13)
+(see Appendix B.4 for the derivation).
+Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$.
+In the special case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection.
+This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods.
+Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting the ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013).
+Fig.
+6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$.
+With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases.
+Compared with the no-selection baseline $(p_{\mathrm{sel}},\pi)=(1,1)$, random selection along the diagonal $p_{\mathrm{sel}}=\pi$ does not reduce the error, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
+Figure 6: Prediction error of sparse regression in the student–teacher model.
+Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right).
+The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$.
+The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection.
+The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
+Sparse regression can exploit a sparse brain-to-feature mapping and is not structurally constrained by ODC.
+Our analysis within the student–teacher framework shows that variable selection, which achieves a high hit rate, can lead to accurate prediction even at small data scales.
+Our error expression in Eq.
+12 applies to a broad class of sparse translators beyond the specific correlation-based filter we study next.
+6 Quantifying hit rate and prediction error in practical sparse regression
+We derived a prediction error expression for a broad class of sparse regression methods in terms of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$.
+We next focus on a simple correlation-based filter as a concrete variable selection method and analyze how it realizes $p_{\mathrm{sel}}$ and $\pi$ under the student–teacher model.
+We first derive a relation between $p_{\mathrm{sel}}$, $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$.
+We then evaluate the validity and practical range of our theoretical analysis by comparing the theoretical prediction errors with those from simulations and fMRI brain-to-image reconstruction data.
+Among the various variable selection methods, we focus on a simple correlation-based filter (akin to SIS (Fan and Lv, 2008)) as a concrete procedure (Fig.
+7A).
+This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data.
+For each output dimension, the filter selects the $d_{\mathrm{sel}}=p_{\mathrm{sel}}d$ input variables based on correlation scores (see Appendix B.5 for details).
+Applying this procedure separately across output dimensions yields a column-wise sparse support, which may differ across outputs.
+Figure 7: Correlation-based variable selection and its hit rate under the student–teacher model.
+(A) Correlation-based filter for variable selection.
+The depth labeled $n$ is the sample axis: each slice represents one training sample.
+For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}=p_{\mathrm{sel}}d$ by absolute correlation, and fit the model using only the selected inputs.
+(B) Hit rate versus selection rate with a correlation-based filter.
+Each curve shows the relationship between the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for different non-zero weight ratios $a$ under the student–teacher model.
+The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$.
+Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
+Under the student–teacher model (Fig.
+5B), we obtain the following asymptotic relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$ as $n,d\to\infty$ with $n/d<\infty$.
+$\pi\approx\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right),$ (14)
+with
+$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\qquad\alpha=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (15)
+where $\Phi$ is the cumulative distribution function of the standard normal distribution.
+A detailed derivation is provided in Appendix B.6.
+This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving Eq.
+14 for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$.
+Fig.
+7B visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$.
+For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$.
+In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
+We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter.
+Under the student–teacher model (Fig.
+5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using Eq.
+12 and Eq.
+14.
+We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions.
+For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under Eq.
+12 and Eq.
+14.
+Fig.
+8A summarizes the theoretical estimates.
+For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data.
+For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller.
+These results suggest that a sparse regression model can leverage the sparse structure of the data to achieve substantial accuracy gains and make accurate predictions even at small data scales when the mapping is sufficiently sparse.
+Figure 8: The theoretical and simulated prediction error.
+(A) The estimated prediction error in the student–teacher model.
+The black line shows the theoretical prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed.
+The colored lines show the theoretical prediction error of sparse regression for different non-zero weight ratios $a$.
+The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$.
+(B) The simulated prediction error under four data generation processes.
+Each plot shows the simulated prediction error $R$ of sparse regression under four different data generation processes: (1) Baseline (the same as the setting in our theory), (2) Gaussian-distributed non-zero weights, (3) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (4) Input noise.
+The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$.
+Each colored line shows the prediction error for a different non-zero weight ratio $a$.
+To evaluate the finite-sample accuracy of our theory and its robustness beyond the modeling assumptions, we conduct simulations under four data generation processes (Fig.
+8B).
+Specifically, we consider: (1) a baseline that exactly matches the assumptions used in our analysis (Fig.
+5B); (2) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (3) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect the correlation structure in neural activity; and (4) input noise representing neural variability and measurement noise (see Appendix B.7 for details).
+Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors.
+For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales.
+This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation.
+These results show that our theoretical error expressions remain accurate in finite samples and are robust to deviations from the modeling assumptions that reflect properties of real data.
+They indicate that our framework provides reliable and informative guidance in practical settings.
+Figure 9: Comparison between theoretical and empirical prediction errors.
+(A) Theoretical prediction error of naive and sparse regression as a function of the data scale $\gamma=n/d$ under the student–teacher model.
+Each curve corresponds to a different non-zero weight ratio $a$, and the noise level is set to $\sigma=5$.
+(B) Empirical prediction errors of naive and sparse translators on the Deeprecon dataset (Shen et al., 2019) as a function of the number of training samples $n\in\{150,300,600,1200\}$.
+Since $d_{\mathrm{out}}$ is very large, these $n$ values correspond to small data scales $\gamma=n/d_{\mathrm{out}}<1$.
+Panels A and B use different error scales, so their absolute magnitudes are not directly comparable; we instead compare the relative errors of naive and sparse regression.
+The empirical errors show only a modest advantage of sparse regression over naive regression.
+Finally, we provide an exploratory link between our theory and the empirical regime of real fMRI decoding on the Deeprecon dataset (Shen et al., 2019).
+We vary the number of training samples $n$ over the range $\{150,300,600,1200\}$ and evaluate the prediction error of both naive and sparse linear regression.
+The dimensionality of the latent features is on the order of $d_{\mathrm{out}}\approx 10^{6}$, so these training sample sizes correspond to small data scales $\gamma=n/d_{\mathrm{out}}<1$, even when accounting for redundancy in the latent representation.
+Existing fMRI datasets likely fall in a relatively high-noise regime.
+Since the trial-to-trial variability for repeated presentations is on the order of five in our noise-to-signal metric (see Appendix B.8), we use $\sigma=5$ as a convenient reference value to illustrate this regime and interpret the real-data comparison qualitatively.
+The true non-zero weight ratio $a$ is unknown, so we evaluate the theoretical error over a range of plausible values of $a$.
+Fig.
+9A shows the theoretical error of naive and sparse regression, and Fig.
+9B shows the empirical prediction errors on the Deeprecon dataset (see Appendix B.9 for details).
+Panels A and B use different error scales, so their absolute magnitudes are not directly comparable; we instead compare the relative errors of naive and sparse regression.
+Within the range of $\gamma$, $\sigma$, and $a$ compatible with this dataset, our analysis predicts that sparse regression yields only a modest reduction in prediction error, while a substantial part of the error remains irreducible due to noise.
+Consistent with this prediction, the empirical results show that sparse regression achieves a modest error reduction compared to naive regression.
+These results suggest that a further reduction of prediction error in existing data will require lowering measurement noise in addition to alleviating ODC.
+7 Discussion
+In this study, we provided a theoretical analysis of translator behavior at small data scales.
+We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluated its impact on the reconstruction pipelines that employ a naive translator.
+We then analyzed sparse linear regression in a student–teacher model and derived an expression for the prediction error in terms of data scale, sparsity, and related parameters.
+Our results suggest that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure.
+These findings imply that sparse regression does not suffer from ODC; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
+In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator.
+This best prediction gives the lowest possible latent-feature error, so no naive translator can outperform it even under noise-free brain activity.
+As a caveat, the smallest latent-feature error need not correspond to the most accurate reconstructed image after applying the generator because latent-feature error and image-space error can diverge.
+Even so, the analysis makes clear that collapse at small data scales imposes a substantial irreducible error on the latent features, which directly limits the reconstruction quality.
+In addition, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al.
+(2025).
+The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment.
+Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
+In Section 5 and Section 6, we analyzed sparse linear regression in a student–teacher model and compared its predictions with simulations and the Deeprecon dataset.
+The theory predicts substantial accuracy gains when the noise level is moderate and the brain-to-feature mapping is sufficiently sparse.
+Current brain-to-image reconstruction data depart from this regime: the estimated noise level is high, and the sparsity of the true mapping is likely limited in practice.
+Under these conditions, the prediction error is dominated by noise, which accounts for the modest gains observed on real data.
+Moreover, strong sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in current reconstruction studies, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely satisfy this assumption.
+Our results do not provide an immediate recipe for large performance improvements on existing datasets, but instead offer a quantitative guideline for future choices of measurement procedures and feature representation design: further progress will require both improved denoising of neural recordings and the construction of AI feature spaces that more closely approximate the brain’s representations, so that strong sparsity can be realized and effectively exploited.
+Although we examined only linear translators, several recent reconstruction methods have adopted nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023; 2024).
+In principle, nonlinear networks can encode additional inductive biases, which may be attractive in data-scarce settings.
+However, simply increasing model capacity does not eliminate ODC as a concern.
+In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid ODC.
+Empirically, Shirakawa et al.
+(2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena.
+A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
+Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction.
+In the latest method of Shen et al.
+(2019), the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses.
+More generally, it has been observed that a generator can recover a visually similar image even under substantial perturbations of the latent features (Onoo et al., 2025; Lee et al., 2025).
+This observation suggests that high image-space accuracy may be achievable even when latent-feature error remains significant.
+More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024).
+These generators can produce photorealistic images even from inaccurate latent features.
+However, this approach carries the risk of introducing information not actually present in the brain’s representation.
+In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity.
+The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
+Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language.
+For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020).
+When solving regression with small datasets, ODC is a critical concern.
+Sparse regression can mitigate ODC; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021).
+Although we do not treat this situation here, understanding how such selection concentration shapes prediction remains an important direction for future work.
+The significance of our study lies in its dual contribution.
+First, we showed that naive linear regression suffers from ODC at small data scales, restricting generalization beyond the training data.
+Second, within a student–teacher framework, we derived a prediction error expression for sparse regression and clarified when accurate prediction is achievable at small data scales.
+These contributions provide a method for diagnosing ODC in existing reconstruction pipelines and offer quantitative guidelines for advancing brain-to-image reconstruction without relying on large-scale data.
+Code availability.
+Code is available at GitHub.
+Acknowledgments
+We thank our laboratory team, especially Ken Shirakawa and Hideki Izumi, for their invaluable feedback and insightful suggestions on the manuscript.
+This work was supported by Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705, JP25H00450 to Y.K., and JP21K17821 to Y.N.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, JPMJCR22P3 to Y.K., and BOOST grant JPMJBS2407 to K.O.), New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.), and Guardian Robot Project, RIKEN.
+Appendix A Supplementary Figures
+Figure A1: Comparison of the best prediction and the brain prediction for subsampled training sets.
+Each panel shows best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ with subsampled training sets of 1200, 600, 300, and 150 samples.
+As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
+Figure A2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023).
 (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$.
@@
 From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$.
-The bottom row shows the latest method reconstructions (Shen et al., 2019).
-The left four columns show natural images, and the right two columns show artificial shape images.
-These results confirm that naive multivariate linear regression suffers severely from output dimension collapse due to small data scales.
-Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction.
-In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions.
-These results underscore that naive regression is incompatible with the fundamental structure of reconstruction data: the vast image space and the limited training data.
-5 Leveraging sparse brain-to-feature structure
-In the previous sections, we demonstrated that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli.
-This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data.
-In contrast, brain-to-feature mapping is known to be sparse, and several studies have reported improved prediction performance by exploiting this sparsity.
-In this section, we demonstrate that sparse linear regression, which leverages this structure, can yield low prediction errors even at small data scales.
-We first show that sparse regression is not subject to output dimension collapse.
-We then formalize a data generation process and estimate the prediction error characterized by a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
-In brain-to-image reconstruction, employing “brain-like” latent features renders the brain-to-feature mapping sparse.
-Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000).
-Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
-Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Figure 5A).
-Figure 5: Sparse brain-to-feature mapping.
-(A) If the brain and the latent features exhibit local or selective activity, their relationship should become sparse mapping.
-(B) Data generation process that reflects the sparse structure of the brain-to-feature mapping.
-The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
-Sparse regression models can effectively exploit the sparse brain-to-feature mapping.
-The model selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters.
-Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)).
-Importantly, these models are not subject to output dimension collapse at small data scales ($n<d_{\mathrm{out}}$).
-Naive regression uses a single shared set of input variables to predict all output dimensions, which implies $\mathrm{rank}(\hat{W})\leq n$.
-In contrast, a sparse model can select a different subset of input variables for each output dimension, allowing $\mathrm{rank}(\hat{W})>n$.
-Conceptually, ODC results from a mismatch between the implicit parameter reduction of the naive model and the data structure.
-When the true mapping is sparse, explicit reduction through variable selection that matches this structure should be preferred.
-Theoretically, sparse regression can achieve accurate prediction with small data scales by leveraging sparsity (Wainwright, 2009a; Donoho, 2006; Fan and Lv, 2008).
-Empirically, reconstruction methods that employ brain-like latent features and sparse regression models have achieved accurate zero-shot reconstruction even with limited training data (Miyawaki et al., 2008; Shen et al., 2019).
-To engage in a deeper discussion, we formalize a data generation process that reflects the sparse structure of the brain-to-feature mapping (Figure 5B).
-Let us assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$ for simplicity.
-Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated as follows:
-$\bm{y}^{(i)}$ $=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)},$ (10)
-where the brain activity $\bm{x}^{(i)}$ is generated by sampling $\bm{x}^{(i)}\sim\mathcal{N}(\bm{0},I_{d})$, the noise $\bm{\epsilon}^{(i)}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{d})$, and the teacher weight matrix $W\in\mathbb{R}^{d\times d}$ is column-wise $s$-sparse:
-$W_{j,k}=\begin{cases}\frac{1}{\sqrt{s}}&j\in S_{k},\\ 0&j\notin S_{k},\end{cases}\quad(j=1,2,\ldots,d),$ (11)
-with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices.
-We denote the non-zero weight ratio by $a=s/d$, reflecting the fraction of input dimensions that contribute to each output dimension.
-The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
-Under this generative setting (Figure 5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$.
-The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support.
-Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected.
-The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
-$R\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\},$ (12)
-with
-$\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},\qquad\tau^{2}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\qquad\rho=\frac{p_{\mathrm{sel}}}{\gamma}$ (13)
-(see A.3 for the derivation).
-Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$.
-In case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection.
-This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods.
-Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013).
-Figure 6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$.
-With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases.
-Along the diagonal $p_{\mathrm{sel}}=\pi$ (random selection), the error does not decrease, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
-Figure 6: Prediction error of sparse regression.
-Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right).
-The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$.
-The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection.
-The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
-Sparse regression reduces prediction error by leveraging the sparse brain-to-feature mapping, a structure that arises when brain-like latent features are employed.
-By formalizing this sparse structure, we have shown that appropriate variable selection can lead to accurate prediction even at small data scales.
-6 Quantifying hit rate and prediction error in practical sparse regression
-In the previous section, we analyzed how the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ jointly determine the prediction error in sparse regression.
-However, the selection algorithm and the data determine these two quantities.
-In this section, we focus on a concrete variable selection procedure and derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$.
-Furthermore, by comparing the prediction errors by our theory with those from simulations, we confirm the accuracy of the analysis and its applicability.
-Among the various variable selection methods, we focus on the simple correlation-based (akin to SIS (Fan and Lv, 2008)) as a concrete method (Figure 7).
-This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data.
-For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
-$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$
-where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations.
-Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $p_{\mathrm{sel}}$ fraction of input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$.
-A separate ridge regression uses only the selected inputs to predict each output dimension.
-Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
-Figure 7: Correlation-based filter.
-The depth labeled $n$ is the sample axis: each slice represents one training sample.
-For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}$ by absolute correlation, and fit the model using only the selected inputs.
-Under the data generation process introduced in Section 5 (Figure 5B), we can derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$.
-We obtain the following relation for large $d$:
-$\pi\approx\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right),$ (14)
-with
-$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\qquad\alpha=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (15)
-where $\Phi$ is the cumulative distribution function of the standard normal distribution.
-A detailed derivation is provided in A.4.
-This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving (14) for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$.
-Figure 8 visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$.
-For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$.
-In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
-Figure 8: Hit rate versus selection rate with a correlation-based filter.
-Each curve shows the estimated prediction error $R$ for different non-zero weight ratios $a$.
-The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$.
-Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
-We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter.
-Under the assumed data generation process (Figure 5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using (12) and (14).
-We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions.
-For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under (12) and (14).
-Figure 9 summarizes the theoretical estimates.
-For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data.
-For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller.
-These results demonstrate that a practical sparse regression model can translate the sparse structure of the data into substantial accuracy gains and that, when the mapping is sufficiently sparse, it can achieve accurate zero-shot prediction even at small data scales.
-Figure 9: The estimated prediction error from theory.
-The black line shows the prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed.
-The colored lines show the prediction error of sparse regression for different non-zero weight ratios $a$.
-The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$.
-To assess the accuracy and applicability of our theory, we conduct simulations under four data generation processes (Figure 10).
-Specifically, we consider: (A) a baseline that exactly matches the assumptions used in our analysis (Figure 5B); (B) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (C) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect correlation structure in neural activity; and (D) input noise representing neural variability and measurement noise (see A.5 for details).
-Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors.
-For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales.
-This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation.
-These results indicate that our framework provides reliable and informative guidance in practical settings.
-Figure 10: The simulated prediction error under four data-generating processes.
-Each plot shows the simulated prediction error $R$ of sparse regression under four different data-generating processes: (A) Baseline (the same as the setting in our theory), (B) Gaussian-distributed non-zero weights, (C) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (D) Input noise.
-The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$.
-Each colored line shows the prediction error for a different non-zero weight ratio $a$.
-7 Discussion
-In this study, we theoretically analyzed how translator models behave under small data scales.
-We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluate its impact on the reconstruction pipeline that employs a naive translator.
-We then mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters.
-Our results show that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure.
-These findings imply that sparse regression does not suffer from output dimension collapse; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
-In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator.
-Although this best prediction minimizes latent-feature error, it does not always yield the most accurate reconstruction.
-Even so, it is evident that the output dimension collapse caused by small data scales prevents the model from accurately capturing the true latent features, thereby compromising the quality of the reconstructed images.
-Moreover, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al.
-(2025).
-The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment.
-Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
-In Section 5 and Section 6, we evaluated sparse linear regression under the idealized assumptions that model real-world data.
-Although real datasets do exhibit a performance boost from sparse regression, the improvement is typically more modest than in our simulations.
-We performed additional simulations under a range of more realistic conditions, yet some discrepancies with actual data are likely to remain.
-First, our simulations cover relatively high SNR regimes.
-Today’s brain data contains more noise (Allen et al., 2022; Hebart et al., 2023), suggesting that the effective noise level is higher than the values used in our simulations.
-Although our theory is applicable regardless of the noise level, obtaining a small prediction error requires sufficiently low noise.
-Second, the latent features used in practice may not fully satisfy the sparsity assumption.
-While our simulations considered several sparsity levels, it is challenging to estimate the degree of sparsity in real data.
-Such sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in practice, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely meet this ideal.
-Improvements in neural recording denoising techniques and the design of AI feature spaces that empirically approximate the brain’s representations will likely enable zero-shot reconstruction at small data scales.
-Although we examined only linear translators, recent reconstruction methods sometimes deploy nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023, 2024).
-For instance, Chen et al.
-(2023) employed Masked Brain Modeling, embedding structural priors of brain activity into the translator in an attempt to improve prediction accuracy.
-Although nonlinear networks can encode such inductive biases, output dimension collapse remains a concern.
-In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid output dimension collapse.
-Empirically, Shirakawa et al.
-(2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena.
-A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
-Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction.
-In the latest method, as described in Section 3, the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses.
-More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024).
-These models enable us to generate photorealistic images even from inaccurate latent features.
-However, this approach carries the risk of introducing information not actually present in the brain’s representation.
-In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity.
-The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
-Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language.
-For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020).
-When solving regression at small data scales, output dimension collapse is a critical concern.
-Sparse regression is effective in avoiding collapse; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021).
-Although we do not treat this situation here, determining how such selection concentration shapes prediction remains an important direction for future work.
-7.1 Conclusion
-The significance of our study lies in its dual contribution: First, we show that naive linear regression suffers from output dimension collapse at small data scales, restricting generalization beyond the training data.
-Second, we demonstrate that sparse linear regression can achieve accurate prediction at small data scales by mathematically characterizing the prediction error.
-These findings provide a practical guideline for advancing image reconstruction without relying on large-scale data.
-8 Declaration of Generative AI and AI-assisted technologies in the writing process
-During the preparation of this work, the authors used ChatGPT (GPT-4o, o3, 5) in order to improve readability and English translation.
-After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
-9 Acknowledgements
-We thank our laboratory team, especially Ken Shirakawa and Hideki Izumi, for their invaluable feedback and insightful suggestions on the manuscript.
-This work was supported by Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705, JP25H00450 to Y.K., and JP21K17821 to Y.N.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, JPMJCR22P3 to Y.K., and BOOST grant JPMJBS2407 to K.O.), New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.), and Guardian Robot Project, RIKEN.
-Appendix A Supplementary Materials
-A.1 Output subspace restriction in kernel linear regression and PLS
-This section shows that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
+Similar results are observed: the brain prediction and the best prediction are close to each other but far from the true feature.
+Appendix B Supplementary Materials
+B.1 Output subspace restriction in kernel linear regression and PLS
+We show that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
+Let $(\bm{x}^{(i)},\bm{y}^{(i)})$ for $i=1,\ldots,n$ be training pairs with $\bm{x}^{(i)}\in\mathbb{R}^{d_{\mathrm{in}}}$ and $\bm{y}^{(i)}\in\mathbb{R}^{d_{\mathrm{out}}}$.
+Let $Y_{\mathrm{tr}}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ be the training output matrix and define $\mathcal{A}:=\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\}$.
 Kernel linear regression.
@@
 $\hat{\bm{y}}_{\mathrm{te}}\;=\;A^{\top}\bm{k}_{\mathrm{te}}\quad\text{with}\quad A\;=\;(K+\lambda I)^{-1}Y_{\mathrm{tr}}.$
-Therefore
+Therefore,
 $\hat{\bm{y}}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}\bm{m}=\sum_{i=1}^{n}m_{i}\,\bm{y}^{(i)},$
-where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$. Thus, every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
+where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$.
+Thus, $\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}$, i.e., every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
 Partial least squares (PLS).
@@
 where $\bm{m}\;=\;T(T^{\top}T)^{-1}t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{n}$.
-Thus every prediction $\hat{\bm{y}}_{\mathrm{te}}$ is a linear combination of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
+Thus, $\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}$.
 This conclusion holds for NIPALS and SIMPLS alike, regardless of the specific construction of $T$ and $t(\cdot)$.
-A.2 The best prediction
+B.2 Derivation of the best prediction for naive linear regression
+Let $\bm{y}_{\mathrm{te}}\in\mathbb{R}^{d_{\mathrm{out}}}$ be a test output and $Y_{\mathrm{tr}}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ be the training output matrix.
 The best prediction in naive multivariate linear regression is derived analytically as follows:
 $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$
-$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
-where $A^{+}$ is the Moore-Penrose pseudo-inverse of $A$.
+$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
+where $A^{\dagger}$ is the Moore-Penrose pseudo-inverse of $A$.
 Proof.
@@
 Thus, the best prediction is
-$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
-A.3 Prediction error for sparse linear regression
+$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
+B.3 ODC analysis in the Deeprecon dataset
+We analyzed output dimension collapse (ODC) on real fMRI data using the Deeprecon dataset (Shen et al., 2019).
+All analyses used subject sub-01 and the visual cortex (VC) ROI.
+Unless stated otherwise, preprocessing, latent feature extraction, and reconstruction followed Shen et al.
+(2019).
+The training set contains 1,200 ImageNet natural images (Deng et al., 2009), and the test set contains 50 natural images from held-out categories and 40 artificial shape images (Shen et al., 2019).
+Our analysis mainly followed the reconstruction pipeline of Shen et al.
+(2019), which employs a Translator–Generator framework in which a linear translator predicts latent features from fMRI and a reconstruction step performs iterative feature inversion to produce an image that matches the target features.
+To analyze ODC under a naive multivariate linear translator, we replaced the sparse translator of Shen et al.
+(2019) with multivariate ridge regression using a fixed penalty $\lambda=1000$.
+Latent feature vectors were standardized using a standard scaler fit on the training outputs, and all projections and errors were computed in this standardized feature space.
+We computed $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ as the orthogonal projection of the standardized true feature $\bm{y}_{\mathrm{te}}$ onto $\mathrm{span}(Y_{\mathrm{tr}})$, implemented as
+$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}=Y_{\mathrm{tr}}^{\top}(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top})^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$ (16)
+To vary the data scale, we subsampled the 1,200 training images to $n\in\{600,300,150\}$ by selecting 4, 2, and 1 images per category (150 categories), respectively, and used one fixed subset for each $n$.
+To visualize target features as images, we used Deep Image Prior (DIP) (Ulyanov et al., 2018) and optimized both the DIP network parameters and its input so that the features extracted from the generated image matched the target feature.
+We used AdamW (learning rate $10^{-3}$) for 800 iterations.
+Prediction error was measured as the normalized squared error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}-\bm{y}\|^{2}$.
+For comparing $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the ridge-based brain prediction $\hat{\bm{y}}_{\mathrm{te}}$, we computed the per-sample ratio $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ and summarized it by the mean and standard deviation across samples.
+In addition, we evaluated the reconstruction method of Ozcelik and VanRullen (Ozcelik and VanRullen, 2023) on the same Deeprecon data.
+We followed their original procedure without modification, and the corresponding results are reported in Fig.
+A2.
+B.4 Derivation of the prediction error for sparse linear regression
 We quantify the prediction error when the sparse regression retains a fraction $p_{\mathrm{sel}}$ of the inputs and captures a fraction $\pi$ of the truly relevant ones.
@@
 This expression gives the prediction error of sparse regression with a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
-A.4 Selection rate-hit rate relation for correlation-based filter
+B.5 Correlation-based variable selection
+To analyze practical sparse regression within our theoretical framework, we considered a simple correlation-based filter (akin to SIS (Fan and Lv, 2008)) as a concrete procedure (see also Fig.
+7A).
+This choice was also motivated by practice: Shen et al.
+(2019) employed correlation-based voxel selection as part of their reconstruction pipeline.
+This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data.
+For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
+$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$
+where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations.
+Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $d_{\mathrm{sel}}$ input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$.
+A separate ridge regression uses only the selected inputs to predict each output dimension.
+Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
+B.6 Derivation of the selection rate and hit rate relation for correlation-based filter
 We relate the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ of the correlation-based filter to characterize the performance of the practical filter.
@@
 For each output coordinate $k\in\{1,\ldots,d\}$, we can rewrite the data generation process as
-$\mathbf{y}_{k}$ $=X_{\mathrm{tr}}W_{:,k}+\mathbf{\epsilon}_{k}$ (31)
-$=\sum_{j=1}^{d}W_{jk}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (32)
-$=\frac{1}{\sqrt{s}}\sum_{j\in\mathcal{S}_{k}}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (33)
-where $W_{:,k}$ is the $k$-th column of the teacher weight matrix $W$, $\mathbf{\epsilon}_{k}\in\mathbb{R}$ is the $k$-th output noise vector, and $\mathcal{S}_{k}=\left\{j\mid W_{jk}\neq 0\right\}$ is the informative index set for the $k$-th output coordinate.
-Within this data generation process, $\mathbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\mathbf{\epsilon}_{k}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{n})$ hold.
+[…]
+[…]
+[…]
+[…]
 The correlation-based filter ranks each input variable $j$ according to its absolute correlation with the output $k$:
@@
 $z_{jk}$ $=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle$ (37)
-$=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (37)
-$=\frac{1}{\sqrt{s}}\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (38)
+[…]
+[…]
 Independence of the Gaussian components yields
-$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (39)
+[…]
 $=\frac{n}{\sqrt{s}}+0+0=\frac{n}{\sqrt{s}},$ (41)
-$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{s}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (41)
+[…]
 $=\frac{2n}{s}+\frac{(s-1)n}{s}+n\sigma^{2}=n(\frac{s+1}{s}+\sigma^{2}).$ (43)
 Similarly, for an uninformative input $j\notin\mathcal{S}_{k}$, we find
-$z_{jk}$ $=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (43)
-$=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (44)
-$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (45)
+[…]
+[…]
+[…]
 $=0+0=0,$ (47)
-$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\sum_{t\in\mathcal{S}_{k}}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (47)
+[…]
 $=n(1+\sigma^{2}).$ (49)
@@
 $=\Phi\!\left(-\frac{u_{k}-\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\Phi\!\left(-\frac{u_{k}+\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (54)
-(51) follows from Berry-Esseen theorem, and (53) follows from
+Eq. 52 follows from Berry-Esseen theorem, and Eq. 54 follows from
 $\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)=\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(1/s\right).$ (55)
@@
 $=2\,\Phi\!\left(-\frac{u_{k}}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (58)
-Combining (53) and (57) with (49) yields the two-equation relation
+Combining Eq. 54 and Eq. 58 with Eq. 50 yields the two-equation relation
 $p_{\mathrm{sel}}$ $=a\!\left[\Phi\!\left(-\tfrac{u_{k}-\Delta}{\sqrt{1+\sigma^{2}}}\right)+\Phi\!\left(-\tfrac{u_{k}+\Delta}{\sqrt{1+\sigma^{2}}}\right)\right]$
@@
 while in the boundary case $a=1$ the relation reduces to $p_{\mathrm{sel}}=\pi$.
-A.5 Detailed Simulation Settings
-We conducted simulations under the following four generative conditions:
+B.7 Simulation study: generative conditions and parameters
+To evaluate the finite-sample accuracy of our theory and its robustness to deviations from the assumptions, we conducted simulations under the following four generative conditions:
   • (A)
@@
 Each non-zero entry in the teacher matrix $W$ is set to $1/\sqrt{s}$, with the remaining entries set to 0.
-The input vector $\bm{x}[{(i)}]$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
+The input vector $\bm{x}^{(i)}$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
   • (B)
@@
   • (D)
-    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)},\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$.
+    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)}$ and $\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$.
 The teacher matrix $W$ has the same binary non-zero structure as in (A).
@@
 The simulations are performed with the following parameters: the dimensionality $d=1000$ and the noise level $\sigma=0.1$.
-Appendix B Supplementary Figures
-Figure 1: Comparison of the best prediction and the brain prediction for subsampled training sets.
-Each panel shows the same analysis as in Figure 4 with subsampled training sets of 1200, 600, 300, and 150 samples.
-Each dot represents one image, and the dotted line indicates where the two errors are equal.
-Blue for natural images, orange for artificial shapes.
-As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
-Figure 2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023).
-(A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$.
-Blue for natural images, orange for artificial shapes.
-The error bars denote the standard deviation across samples.
-Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$.
-Each dot represents one image, and the dotted line indicates where the two errors are equal.
-(B) Representative reconstructions generated from the best predictions and brain predictions.
-From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$.
-Similar results are observed: the brain prediction and best prediction are close to each other but far from the true feature.
+For ridge regression, we used the ridge penalty that is theoretically optimal in the baseline setting, $\lambda=d/\tau^{2}$, in all simulation conditions (A–D).
+B.8 Noise level estimation in the Deeprecon dataset
+To set the noise level used in our theoretical comparisons with real fMRI results, we estimated the trial-to-trial noise variance from repeated presentations of identical stimuli in the Deeprecon dataset (Shen et al., 2019).
+We used an ANOVA-like variance decomposition to separately estimate the stimulus-driven variance and the within-stimulus trial noise.
+For each voxel, let $x_{i,r}$ denote the preprocessed fMRI response to stimulus $i\in\{1,\ldots,s\}$ at repetition $r\in\{1,\ldots,t\}$. We defined the per-stimulus mean and the grand mean as
+$\bar{x}_{i\cdot}=\frac{1}{t}\sum_{r=1}^{t}x_{i,r},\qquad\bar{x}_{\cdot\cdot}=\frac{1}{st}\sum_{i=1}^{s}\sum_{r=1}^{t}x_{i,r}.$ (63)
+We then computed the between-stimulus and within-stimulus sums of squares,
+$\mathrm{SS}_{\mathrm{between}}$ $=t\sum_{i=1}^{s}(\bar{x}_{i\cdot}-\bar{x}_{\cdot\cdot})^{2},$ (64)
+$\mathrm{SS}_{\mathrm{within}}$ $=\sum_{i=1}^{s}\sum_{r=1}^{t}(x_{i,r}-\bar{x}_{i\cdot})^{2},$ (65)
+and their corresponding mean squares,
+$\mathrm{MS}_{\mathrm{between}}=\frac{\mathrm{SS}_{\mathrm{between}}}{s-1},\qquad\mathrm{MS}_{\mathrm{within}}=\frac{\mathrm{SS}_{\mathrm{within}}}{s(t-1)}.$ (66)
+We modeled repeated responses by a random-effects decomposition
+$x_{i,r}=\mu+z_{i}+\varepsilon_{i,r},$ (67)
+where $z_{i}\sim\mathcal{N}(0,\sigma_{z}^{2})$ captures stimulus-driven variability across stimuli and $\varepsilon_{i,r}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$ captures trial-to-trial noise.
+Under this model, the expected mean squares satisfy $\mathbb{E}[\mathrm{MS}_{\mathrm{within}}]=\sigma_{\varepsilon}^{2}$ and $\mathbb{E}[\mathrm{MS}_{\mathrm{between}}]=\sigma_{\varepsilon}^{2}+t\sigma_{z}^{2}$.
+We therefore estimated
+$\widehat{\sigma}_{\varepsilon}^{2}=\mathrm{MS}_{\mathrm{within}},\qquad\widehat{\sigma}_{z}^{2}=\frac{\mathrm{MS}_{\mathrm{between}}-\mathrm{MS}_{\mathrm{within}}}{t}.$ (68)
+We computed these estimates voxel-wise and then averaged $\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}$ over voxels with $\widehat{\sigma}_{z}^{2}>0$.
+Using ImageNetTraining (1200 stimuli, 5 repetitions), we obtained $\sigma_{\varepsilon}/\sigma_{z}=\sqrt{\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}}\approx 5.7$.
+Using ImageNetTest (50 stimuli, 24 repetitions), we obtained $\sigma_{\varepsilon}/\sigma_{z}=\sqrt{\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}}\approx 5.2$.
+Overall, these estimates suggest that the noise level $\sigma_{\varepsilon}/\sigma_{z}$ was on the order of $5$ and was relatively consistent across the two datasets.
+B.9 Naive vs. sparse translator comparison in the Deeprecon dataset
+We compared naive and sparse translators on the Deeprecon dataset (Shen et al., 2019) using the same subject (sub-01), VC ROI, preprocessing, and category-balanced subsampling protocol as in the ODC analysis.
+We evaluated prediction error on the 50 natural test images (held-out categories).
+The naive translator was multivariate ridge regression with a fixed penalty $\lambda=1000$.
+The sparse translator was a correlation-based filter plus ridge regression: we used a fixed ridge penalty $\lambda=100$ and set the number of selected voxels per output dimension to $d_{\mathrm{sel}}\in\{63,125,250,500\}$ for $n\in\{150,300,600,1200\}$, respectively.
+Prediction error was measured as the normalized squared error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}-\bm{y}\|^{2}$ and averaged across the natural test stimuli.